Calculate Karl Pearson’s coefficient of correlation from the following data taking deviations from Arithmetic mean and interpret the result.



1
Expert's answer
2021-09-14T06:05:24-0400

Karl Pearson's correlation coefficient (r) formula is given by -


"r = \frac{\sum(X-\bar{X})(Y-\bar{Y})}{\sqrt{(\sum(X-\bar{X})^2)\ *\ ({\sum(Y-\bar{Y})^2)}}}"